OpenStudy (jareeek):

Will fan and medal :D The Baldwin family walked four over five of a mile in four over nine of an hour. What is their unit rate in miles per hour? four-ninths over four-fifths equals four over nine mile per hour four-fifths over four-ninths equals one and four-fifths miles per hour four-ninths over four-fifths equals one and four-fifths miles per hour four-fifths over four-ninths equals four over nine mile per hour

OpenStudy (jareeek):

B?

OpenStudy (jareeek):

@Tom_Boy_Rebel

OpenStudy (anonymous):

Let me see...

OpenStudy (anonymous):

Honestly... I am not sure.... It might be C.

OpenStudy (jareeek):

ik its not A cause i got that wrong

OpenStudy (anonymous):

@TheSmartOne @Whitemonsterbunny17 Can you please help @jareeek

OpenStudy (anonymous):

Sorry about that...

OpenStudy (jareeek):

i think its B but im not sure @TheSmartOne

OpenStudy (jareeek):

@pooja195

TheSmartOne (thesmartone):

\(\Large \frac{4}{5} \times \frac{9}{4}=?\)

TheSmartOne (thesmartone):

so correcto :)

OpenStudy (jareeek):

:D

OpenStudy (jareeek):

so B

TheSmartOne (thesmartone):

good job! :)

OpenStudy (jareeek):

Thanks

TheSmartOne (thesmartone):

Anytime! :D